OpenStudy (carniel):

Divide: 36xy - 39y 3y

sam (.sam.):

12x - 13

OpenStudy (anonymous):

\[\frac{36xy - 39y}{3y} = \frac{36xy}{3y}- \frac{39y}{3y}= 12x-13\]